Ellie Mosk, CEO of X-Space Industries, decided to expand the company’s product offering beyond the core model rocket business. After investigation, she decided to set up a separate division to design and manufacture products for the drone market. Several companies were interested in having X-Space develop these drones, and financial results, to date, have been encouraging. Revenue was $4 million, gross margins have been running about 40%, and the customer sales and support costs were $1 million. However, there is a growing concern that some customers require a disproportionate share of the sales and support resources, and the true profitability of the customers is unknown. Data were collected to support an analysis of customer profitability:

Activity Cost Driver Total Cost
Sales visits Sales visit days $ 480,000
Product modifications Number of modifications 266,000
Phone calls Number of minutes 92,600
E-mail/electronic communications Number of communications 168,000
$ 1,006,600
Customer Revenue Gross Profit Visit Days Modifications Phone Minutes Electronic Communications
A $ 394,000 $ 144,000 15 15 1,090 625
B 494,000 194,000 25 15 1,180 875
C 594,000 224,000 40 40 1,430 1,060
D 1,060,000 414,000 90 60 1,780 2,060
E 1,440,000 584,000 100 70 2,180 2,310
Totals $ 3,982,000 $ 1,560,000 270 200 7,660 6,930

Required:

1. Management felt the easiest way to allocate the sales and support costs was based on the total revenue. Using total revenue as the allocation base, determine the profitability of each of the five customers.

2. Management felt that because the data revealed some customers require a disproportionate share of sales and support resources, activity-based costing should be used to determine customer profitability. Use ABC to prepare a customer profitability analysis.
Answer #1

(1): Here the total support costs of $1,006,600 will be allocated on the basis of total revenue. So the amount to be allocated per $ of revenue will be = 1,006,600/3,982,000 = 0.252788

So the amount allocated to customer A will be = $394,000*0.252788 = 99,598.29. In the same manner we can allocate the total support cost to other customers by multiplying their revenue with 0.252788

i ii iii iv = ii - iii
Customer Revenue Gross profit Support costs Profit = gross profit - support costs
A                                    394,000           144,000                    99,598.29                                                           44,401.71
B                                    494,000           194,000                  124,877.05                                                           69,122.95
C                                       594,000           224,000                  150,155.80                                                           73,844.20
D                                 1,060,000           414,000                  267,954.80                                                         146,045.20
E                                 1,440,000           584,000                  364,014.06                                                         219,985.94
Total                                 3,982,000        1,560,000              1,006,600.00                                                         553,400.00

(2): Here we will compute the cost base for each activity. Take the activity of sales visits. Here the cost is $480,000 and the cost driver is sales visit days. Total sales visit days = 270. Thus cost per visit = 480,000/270 = 1,777.78. Thus cost of $480,000 that will be allocated to customer A will be = 480,000/270*15 = $26,666.67. In the same manner we will use the total of cost driver activities and compute cost for each customer.

i ii iii iv v vi vii = ii-iii-iv-v-vi
Customer Revenue Gross profit Sales visit costs Product modification cost Phone calls cost e-mail costs Profit = gross profit - sum of all support costs
A                                    394,000           144,000                    26,666.67                           19,950                  13,176.76                  15,151.52                                       69,055.06
B                                    494,000           194,000                    44,444.44                           19,950                  14,264.75                  21,212.12                                       94,128.68
C                                       594,000           224,000                    71,111.11                           53,200                  17,286.95                  25,696.97                                       56,704.97
D                                 1,060,000           414,000                  160,000.00                           79,800                  21,518.02                  49,939.39                                     102,742.59
E                                 1,440,000           584,000                  177,777.78                           93,100                  26,353.52                  56,000.00                                     230,768.70
Total                                 3,982,000        1,560,000                        480,000                         266,000                        92,600                     168,000                                           553,400